OpenStudy (fanduekisses):

PLEASE HELP ;'( How Do I prove cot(x-pi/2)=-tanx

OpenStudy (fanduekisses):

\[\cot \left( x -\frac{ \pi }{2}\right)=-tanx\]

OpenStudy (anonymous):

How have you defined \(\cot(x)\)?

OpenStudy (anonymous):

What identity do you get to assume?

OpenStudy (fanduekisses):

1/sec ?

OpenStudy (anonymous):

I define \[ \cot(x) = \tan(\pi/2-x) \]

OpenStudy (fanduekisses):

which one is that?

OpenStudy (anonymous):

Because \(\text {co}\) for any trig function means \(f(\pi/2-x)\).

OpenStudy (anonymous):

This is called the co-function identity.

OpenStudy (fanduekisses):

ohhh

OpenStudy (anonymous):

So \(\cos(x) = \sin(\pi/2-x)\).

OpenStudy (anonymous):

And \(\csc(x) = \sec(\pi/2-x)\)

OpenStudy (anonymous):

Also \[ \text{cvs}(x) = \text{ver}(\pi/2-x) \]

OpenStudy (fanduekisses):

oh ok ^_^ thanks

OpenStudy (anonymous):

However, this doesn't fully complete the proof.

OpenStudy (anonymous):

y = cot (x - Pi/2) = cos (Pi/2 - x)/sin (x - Pi/2) Numerator: cos (x - Pi/2) = cos ( Pi/2 - x) = sin x Denominator: sin (x - Pi/2) = -sin (Pi/2 - x) = -cos x Finally: y = sin x/(-cos x) = -tan x